OpenStudy (anonymous):

The y-intercept of the graph of 3x-2y=8

OpenStudy (anonymous):

3x-2y=8 REMEMBER: y-intercept occurs when x=0 3x - 2y = 8 3(0) - 2y = 8 0 - 2y = 8 -2y = 8 --- ---- -2 -2 y = -4 SO: y-intercept is -4 at point (0,-4)

OpenStudy (anonymous):

How did you get the 0?

OpenStudy (anonymous):

to find the y-intercept, X = 0

OpenStudy (anonymous):

and we found out that Y = -4

OpenStudy (anonymous):

X = 0, Y = -4 (x, y,) (0, -4)

OpenStudy (anonymous):

So the x=0 is a standard? So you could use it in other slope problems like this one?

OpenStudy (anonymous):

To find the X-INTERCEPT, let Y=0. To find the Y-INTERCEPT, let X=0.

OpenStudy (anonymous):

Okay. I get it. Thanks!

OpenStudy (anonymous):

yw